With a 10% increase in the price of a dairy products, the number of units supply rises from 

20 to 25. Determine the price elasticity of supply. State whether the supply for the product

is elastic or inelastic.



1
Expert's answer
2021-05-04T07:17:12-0400

Price elasticity of supply"=\frac{Percentage\ change\ in\ quantity}{Percentage\ change\ in\ price}"


"\%\ Change\ in\ quantity=\ \frac{Q_2-Q_1}{(Q_2+Q_1)/2}\times100"


In this case,"\ Q_2=25,\ Q_1=20."


Therefore, "\frac{25-20}{(25+20)/2}\times100"


"=\ 22.22"


"\%\ Change\ in\ price=10\%"


Price elasticity of supply"=\frac{22.22}{10}"


"=2.22"


The supply for the product is elastic because the elasticity is greater than 1 (>1) which indicates high responsiveness to changes in price.